OpenStudy (anonymous):

The following function defines a recursive sequence. f(0) = -2 f(1) = 8 f(n) = -4•f(n -1) - 3•f(n - 2); for n > 1 Which of the following sequences is defined by this recursive function? -2, 8, -26, -80, … -2, 8, -26, 80, … -2, 12, -44, 180, … -2, -12, -44, -180, …

OpenStudy (anonymous):

well you know it is one of the top two since f1 and f2 are given... what happens when you look for f(2)? with F(0) being f(n-2) and f(1) being f(n-1)

OpenStudy (anonymous):

so f(2) = (4*-2) - 3(8)

OpenStudy (anonymous):

ok well obviously its -26... they both are... so next one

OpenStudy (anonymous):

First one?

OpenStudy (anonymous):

f(3) = 4(8)-3(-26) what is it?

OpenStudy (anonymous):

I'm not sure how to solve it

OpenStudy (anonymous):

you know what i plugged in the wrong way... i said it write them typed it wrong . -4*8 -3*(-2)=-26

OpenStudy (anonymous):

Oh okay

OpenStudy (anonymous):

ok the idea with recursion is that you use the answer theat you obtain to get the next in the series

OpenStudy (anonymous):

so of course.. mess one up... mess all the ones after it up .. lol

OpenStudy (anonymous):

so you have the first two values in the series... f(0) and f(1) .... you are looking for f(2) f(3) .... to fin the f(2) say n=2 so f(n-1) = f(1) NOT 1... and f(n-2)=f(0) NOT the value 0

OpenStudy (anonymous):

So the next number will be a positive?

OpenStudy (anonymous):

so we know f(0)=-2 f(1)=8 and now that f(2)=-26 and F(3) makes n=3 so f(n-1)=-26 and f(n-2)=8 we get f(3) = -4*(-26) -3(8) yes it will be positive because the +104> -24 so 80

OpenStudy (anonymous):

sorry took so long had a call